a cabin is traveling with the current went 48 miles in 2 hours. against the current, it took 3 hours to travel the same distance. find the rate of the current
Answer by mananth(16949)

let cabin speed be x mph
current speed = y mph
..
speed downstream = x+y
distance = speed * time
48 = (x+y)*2
/2
24=x+y...................1
...
speed upstream = x-y
3*(x-y)=48
/3
x-y=12...................2add 1&2
2x= 24+12
2x=36
/2
x=18 mph
..
x+y=24
x=18
y=6 mph speed of current